OpenStudy (anonymous):

22% of what number is 33?

hero (hero):

I thought that maybe you would have this figured out by now :(

hero (hero):

Actually, ZhangYan, you kinda have it backwards a little bit

hero (hero):

\[\frac{33}{x}=\frac{22}{100}\]

OpenStudy (anonymous):

no i dont think thts rite

hero (hero):

22x = 3300

hero (hero):

x = 3300/22 x = 150

OpenStudy (anonymous):

22% of 33 = 7.26

OpenStudy (anonymous):

The answer is 150

OpenStudy (anonymous):

7.26

OpenStudy (anonymous):

lol my mistake yes hero is correct. the answer is 150

OpenStudy (anonymous):

its ok thx hero once again :P

OpenStudy (anonymous):

7.26:)

hero (hero):

Musicman, you're confused if you think that is correct

hero (hero):

You haven't even shown any steps

OpenStudy (anonymous):

is 26 of 160 = 41.6 ?

hero (hero):

There's something missing in your statement

OpenStudy (anonymous):

What is 26% of 160? i got 41.6

hero (hero):

That is correct, but I'm more interested in knowing how you got that

OpenStudy (anonymous):

i worked my magic haha no i did my work its pretty easy now i went bck and viewed the steps u showed

hero (hero):

Okay, good

OpenStudy (anonymous):

:)